OpenStudy (anonymous):

Please help simplify this expression and assume all variables are real numbers 5 to the square root of 32(x+4)^5

OpenStudy (anonymous):

it must be 5th root of 32(x+4)^5. the answer then is 2(x+4)

OpenStudy (anonymous):

\[\sqrt[5]{32(x+4)^5}=2(x+4)\]

OpenStudy (anonymous):

Yes you are right. Sorry I don't have the symbols to use

OpenStudy (anonymous):

can you show me or explain how you got that answer please?

OpenStudy (anonymous):

\[\sqrt[5]{32(x+4)^5}=\sqrt[5]{2^5(x+4)^5}=2(x+4)\]

OpenStudy (anonymous):

did u get it? u 2 to the power 5 and (x+4) to the power 5 under the 5th root.

OpenStudy (anonymous):

I think I got it. The power of 5 under the 5th root cancel each other out? Is that right?

OpenStudy (anonymous):

yes..you got it :)

OpenStudy (anonymous):

Cool. Thank you sooo much. I appreciate your help.